Imagine you win $25 million playing the lottery. You, in all your glory, decide that you will donate
the money to revitalizing the US economy which is—at the time of your win—in a recession. You
can choose to give your money to one of two groups. The first group—Group Albatross—has a
marginal propensity to consume of 0.8. The other group—Group Bonobo—has a marginal
propensity to consume of 0.3. If your goal is to jumpstart the economy, which group are you going
to donate your money to? Why?

2) Lucinda received a loan of $100 for a period of one year at an interest rate of 10% from her rich
friend, Kathy. Over the period of a year, the general price level rose by 8%. Who is happy on account
of the price level increase, Kathy or Lucinda? Why?

3) What is the central argument of the AD-AS model? What assumptions is that argument based on?
Answer #1

Q1. For the first group i.e., Group Albatross, Marginal propensity to consume = 0.8

Therefore, spending multiplier = 1/(1-MPC) = 1/(1-0.8) = 1/0.2 = 5

For the second group i.e., Group Bonobo, the marginal propensity to consume = 0.3

Therefore, spending multiplier = 1/(1-MPC) = 1/(1-0.3) = 1/0.7 = 1.43

Therefore, if the goal is to jumpstart the economy, the boost/increase in aggregate demand should be larger. The spending multiplier denotes the number of times the aggregate demand increases compared to the initial spending. Therefore, for given initial spending, the aggregate demand increases by 5 times the initial spending by the Group Albatross and by 1.43 times the initial spending by the group Bonobo

Hence, the money should be donated to the group Albatross